# Simulation of Fermionic circuits using Majorana Propagation

### v1

@@ Abstract
We introduce Majorana Propagation, an algorithmic framework for the classical simulation of Fermionic circuits. Inspired by Pauli Propagation, Majorana Propagation operates by applying successive truncations throughout the Heisenberg evolution of the observable. We identify monomial length as an effective truncation strategy for typical, unstructured circuits by proving that high-length Majorana monomials are exponentially unlikely to contribute to expectation values and the backflow of high-length monomials to lower-length monomials is quadratically suppressed. Majorana Propagation can be used either independently, or in conjunction with quantum hardware, to simulate Fermionic systems relevant to quantum chemistry and condensed matter. We exemplify this by using Majorana Propagation to find circuits that approximate ground states for strongly correlated systems of up to 52 Fermionic modes. Our results indicate that Majorana Propagation is orders of magnitude faster and more accurate than state-of-the-art tensor-network-based circuit simulators.
I Introduction
@@
In this context, Pauli Propagation (PP) has recently emerged as a new promising method for spin systems [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. By leveraging typical features in the dynamics of Pauli strings evolving through a circuit, PP disregards terms that are expected to have a negligible contribution to the expectation value one is simulating, resulting in guaranteed accuracy for almost all circuits within a vast class. PP thus presents itself as a very natural way to simulate expectation values of spin circuits. Despite its recent conception, PP can already compete with well-established methods, such as tensor networks and neural network states. Furthermore, it allows for fast circuit optimization in variational scenarios thanks to surrogate simulation [6], and can be readily interfaced with quantum hardware [12, 15], as it operates at the circuit level.
However, the applicability of PP to other many-body systems, such as Fermionic ones, remains an open question. In particular, many quantum systems are mathematically described by algebras that do not have an equivalent to Pauli weight locality, a quantity central to PP. While these systems are mappable to qubit space, the resulting Pauli operators are typically highly non-local, severely limiting the practicality of PP.
Here we present Majorana Propagation (MP), a PP-inspired classical simulation algorithm that can be naturally used for Fermionic systems. Differently from PP, MP does not operate at the level of Pauli weight; in fact, it does not even assume the decomposability of the algebra describing the system into tensor products of sub-system operators. Instead, MP leverages the typical behaviour of monomials spanning operator bases. Crucially, the Fermionic basis is effectively built ‘on the fly’ and so is not restricted to simulating free Fermion models but rather can be used to simulate systems living in exponentially large Lie algebras. Moreover, unlike tensor network methods or near-Clifford methods, our approach is not intrinsically targeting low entanglement or magic growth scenarios
A particularly promising application of MP is for optimizing circuits to prepare approximate ground states. Established quantum algorithms for energetic structure calculations, including near-term [16, 17, 18, 19, 20, 21, 22, 23, 24, 25] and fault-tolerant [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36] approaches, need to be initialized with good reference states, as their performance depends on the overlap between the reference and the ground state. How to find these remains an open question [37, 38], and heuristic variational ground state approximation methods are likely to play an important role.
@@
Beyond ground states, MP may find use both as a stand-alone classical algorithm, and as a tool to be used in conjunction with quantum hardware, to study Fermionic systems. Potential applications range from simulating the dynamics of Fermionic systems [12, 15] to tackling classification problems with quantum data [50, 52].
II Background
An $N$-mode Fermionic system in second quantization can be described in terms of $N$ creation operators $\{a_{i}^{\dagger}\}_{i=1}^{N}$ and annihilation operators $\{a_{i}\}_{i=1}^{N}$ that satisfy the canonical anticommutation relations $\{a_{i},a_{j}\}=\{a_{i}^{\dagger},a_{j}^{\dagger}\}=0$, and $\quad\{a_{i}^{\dagger},a_{j}\}=\delta_{ij}$. These operators act on the Fermionic Fock space $\mathcal{F}(\mathbb{C}^{N})\cong\oplus^{N}_{k=0}\wedge^{k}\mathbb{C}^{N}$, which is a $2^{N}$-dimensional Hilbert space spanned by the so-called Fock basis. In this space, the Fermionic vacuum |$\ket{\text{vac}_{\text{f}}}$ the unique vector such that $a_{j}\ket{\text{vac}_{\text{f}}}=0$ for all $j=1,\dots,N$. The remaining Fock basis elements can be constructed by considering all possible combinations of occupation numbers $n_{j}\in\{0,1\}$ as |$\ket{n_{1}n_{2}\dots n_{N}}\coloneqq\prod_{j=1}^{N}(a_{j}^{\dagger})^{n_{j}}\ket{\text{vac}_{\text{f}}}$.
It is also common to define the Fermionic space with so-called Majorana operators $\{m_{k}\}_{k=1}^{2N}$ as $m_{2j-1}\coloneqq a_{j}^{\dagger}+a_{j}$ and $m_{2j}\coloneqq\mathrm{i}(a_{j}^{\dagger}-a_{j})$. Such operators obey many convenient properties, such as being unitary and self-adjoint, and satisfy simpler anti-commutation relations:
@@ II Background
The imaginary unit $\mathrm{i}$ is included to ensure that $M_{\boldsymbol{b}}$ is Hermitian [^1] The $\mathrm{i}$ is necessary whenever inverting the product sequence involves an odd number of permutations. That is, the value of $r_{\boldsymbol{b}}$ depends on the length $w=\norm{\boldsymbol{b}}_{1}$ of the monomial. As shown in Appendix A, if either $w$ or $w-1$ is a multiple of 4, $r=0$. Otherwise, $r=1$.. The $4^{N}$ operators $\{M_{\boldsymbol{b}}\}_{\boldsymbol{b}}$ form a basis of the space of linear operators in the Hilbert space, thus any operator can be uniquely decomposed as a linear combination of them. We will refer to these operators as Majorana monomials. The length of a Majorana monomial, defined as the 1-norm of vector $\boldsymbol{b}$, $\norm{\boldsymbol{b}}_{1}=\sum_{i=1}^{2N}b_{i}$, will play a crucial role in this work. We will use the shorthand $w$-monomial to refer to length-$w$ Majorana monomials.
Majorana monomials share some algebraic similarities with Pauli strings: They are traceless (except for the length-0 monomial, i.e., the identity operator). Any two of them either commute or anti-commute, and they square up to identity $M_{\boldsymbol{b}}^{2}=\mathbb{I}$. However, unlike Pauli strings, they are not decomposable as a tensor product of $N$ single-mode operators, as Majorana operators are non-local. In fact, under fermion-to-qubit transformations, Majorana monomials are generally mapped to non-local Pauli strings [39], with Pauli weights spanning a wide range, even for fixed Majorana monomial length. This hinders the implementation of PP for Fermionic systems, as Pauli weight does not capture typical operator dynamics.
III Majorana monomial propagation
@@
$\langle H\rangle_{\boldsymbol{\theta}}={\rm Tr}\left[U(\boldsymbol{\theta})\varrho U(\boldsymbol{\theta})^{\dagger}H\right]$ (3)
given a unitary Fermionic circuit $U(\boldsymbol{\theta})$, an initial state $\varrho$, and an observable $H$. The initial state is typically a Fock basis state, $\varrho=\outerproduct{n_{1}n_{2}\dots n_{N}}{n_{1}n_{2}\dots n_{N}}$ |. We further suppose the Fermionic circuit,
$U(\boldsymbol{\theta})=\prod_{j=1}^{L}e^{-\mathrm{i}\theta_{j}M_{{\boldsymbol{b}_{j}}}/2}\,,$ (4)
@@
If the rotation angles for the circuit one wishes to simulate are known in advance then the sine and cosine terms take numerical values and the $c_{\boldsymbol{b}}\left(\boldsymbol{\theta}\right)$ coefficients can be stored numerically. However, if the values of $\boldsymbol{\theta}$ are not known in advance, or one is interested in computing the expectation value in Eq. (3) for a wide range of $\boldsymbol{\theta}$ values, one can instead symbolically represent the coefficients $c_{\boldsymbol{b}}\left(\boldsymbol{\theta}\right)$. Such surrogate models have a substantial time and memory overhead, but after this pre-processing step, enable rapid re-evaluation of the expectation value [6, 8, 12].
@@
[…]
@@
The second truncation strategy, whereby one aims to cut terms from the sum in Eq. (6) where $\text{Tr}[\varrho M_{\boldsymbol{b}}]$ is likely small, is more subtle. While potentially there are a number of different strategies one could take here, we will make use of the observation that in order for a term $\text{Tr}[\varrho M_{\boldsymbol{b}}]$ to be non-zero all the Majorana operators in the monomial $M_{\boldsymbol{b}}$ are paired. That is, $\text{Tr}[\varrho M_{\boldsymbol{b}}]\neq 0$ if and only if for every even Majorana operator $m_{2i}$ in the Majorana operator $m_{2i+1}$ is also in the monomial (this can be seen immediately from the definition of the Majorana operator, Eq. (16), and noting that $\langle\phi|a_{i}|\phi\rangle=\langle\phi|a_{i}^{\dagger}|\phi\rangle=0$ for any Fock basis state $|\phi\rangle$). Next we observe that in general the higher the length $w$ of the Majorana monomial (for $w<N$) the lower the probability that the monomial is paired. This motivates a length truncation strategy. At each step, after each gate is applied, all terms corresponding to Majorana monomials with length above a certain threshold are to be discarded. This idea is depicted in Fig. 1 and analyzed in more detail in the next section.
IV Length Truncation

### v3

@@ Abstract
We introduce Majorana Propagation, an algorithmic framework for the classical simulation of Fermionic circuits. Inspired by Pauli Propagation, Majorana Propagation operates by applying successive truncations throughout the Heisenberg evolution of the observable. We identify monomial length as an effective truncation strategy for typical, unstructured circuits by proving that high-length Majorana monomials are exponentially unlikely to contribute to expectation values and the backflow of high-length monomials to lower-length monomials is quadratically suppressed. We provide performance guarantees by proving analytically that approximation errors decrease exponentially with the truncation threshold and that only polynomial resources are required to compute the expectation value of observables up to a fixed error for an ensemble of circuits relevant to quantum chemistry. Majorana Propagation can be used either independently, or in conjunction with quantum hardware, to simulate Fermionic systems relevant to quantum chemistry and condensed matter. We exemplify this by using Majorana Propagation to find circuits that approximate ground states for strongly correlated systems of up to 52 Fermionic modes. Our results indicate that Majorana Propagation is orders of magnitude faster and more accurate than state-of-the-art tensor-network-based circuit simulators.
@@
In this context, Pauli Propagation (PP) has recently emerged as a new promising method for spin systems [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 14]. By leveraging typical features in the dynamics of Pauli strings evolving through a circuit, PP disregards terms that are expected to have a negligible contribution to the expectation value one is simulating, resulting in guaranteed accuracy for almost all circuits within a vast class. PP thus presents itself as a very natural way to simulate expectation values of spin circuits. Despite its recent conception, PP can already compete with well-established methods, such as tensor networks and neural network states. Furthermore, it allows for fast circuit optimization in variational scenarios thanks to surrogate simulation [6], and can be readily interfaced with quantum hardware [12, 15], as it operates at the circuit level.
However, the applicability of PP to other many-body systems, such as Fermionic ones, remains an open question. In particular, many quantum systems are mathematically described by algebras that do not have an equivalent to Pauli weight locality, a quantity central to guarantees for PP for typical circuits. While these systems are mappable to qubit space, the resulting Pauli operators are typically highly non-local.
Here we present Majorana Propagation (MP), a PP-inspired classical simulation algorithm that can be naturally used for Fermionic systems. Differently from PP, MP does not operate at the level of Pauli weight; in fact, it does not even assume the decomposability of the algebra describing the system into tensor products of sub-system operators. Instead, MP leverages the typical behaviour of monomials spanning operator bases. We use this observation to prove that Majorana Propagation can efficiently simulate typical Fermionic circuits. Crucially, the Fermionic basis is effectively built ‘on the fly’ and so is not restricted to simulating free Fermion models but rather can be used to simulate systems living in exponentially large Lie algebras. Moreover, unlike tensor network methods or near-Clifford methods, our approach is not intrinsically targeting low entanglement or magic growth scenarios.
A particularly promising application of MP is for optimizing circuits to prepare approximate ground states. Established quantum algorithms for energetic structure calculations, including near-term [17, 18, 19, 20, 21, 22, 23, 24, 25, 26] and fault-tolerant [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38] approaches, need to be initialized with good reference states, as their performance depends on the overlap between the reference and the ground state. How to find these remains an open question [39, 40], and heuristic variational ground state approximation methods are likely to play an important role.
@@
Beyond ground states, MP may find use both as a stand-alone classical algorithm, and as a tool to be used with quantum hardware, to study Fermionic systems. Potential applications range from simulating the dynamics of Fermionic systems [12, 15] to tackling classification problems with quantum data [52, 54].
Background.
[…]
It is also common to define the Fermionic operator space with so-called Majorana operators $\{m_{k}\}_{k=1}^{2N}$ as $m_{2j-1}\coloneqq a_{j}^{\dagger}+a_{j}$ and $m_{2j}\coloneqq\mathrm{i}(a_{j}^{\dagger}-a_{j})$. Such operators obey many convenient properties, such as being unitary and self-adjoint, and satisfy simpler anti-commutation relations:
@@
The imaginary unit $\mathrm{i}$ is included to ensure that $M_{\boldsymbol{b}}$ is Hermitian [^1] The $\mathrm{i}$ is necessary whenever inverting the product sequence involves an odd number of permutations. That is, the value of $r_{\boldsymbol{b}}$ depends on the length $w=\norm{\boldsymbol{b}}_{1}$ of the monomial. As shown in Appendix A, if either $w$ or w1w-1 is a multiple of 4, $r=0$. Otherwise, $r=1$.. The $4^{N}$ operators $\{M_{\boldsymbol{b}}\}_{\boldsymbol{b}}$ form a basis of the space of linear operators in the Hilbert space, thus any operator can be uniquely decomposed as a linear combination of them. We will refer to these operators as Majorana monomials. The length of a Majorana monomial, defined as the 1-norm of vector $\boldsymbol{b}$, $\norm{\boldsymbol{b}}_{1}=\sum_{i=1}^{2N}b_{i}$, will play a crucial role in this work. We will use the shorthand $w$-monomial to refer to length-$w$ Majorana monomials.
Majorana monomials share some algebraic similarities with Pauli strings: They are traceless (except for the length-0 monomial, i.e., the identity operator). Any two of them either commute or anti-commute, and they square up to identity $M_{\boldsymbol{b}}^{2}=\mathbb{I}$. However, unlike Pauli strings, they are not decomposable as a tensor product of $N$ single-mode operators, as Majorana operators are non-local. In fact, under fermion-to-qubit transformations, Majorana monomials are generally mapped to non-local Pauli strings [41], with Pauli weights spanning a wide range, even for fixed Majorana monomial length. It follows that previous guarantees for PP with weight truncation do not apply to Fermionic systems, as Pauli weight does not capture typical operator dynamics.
Majorana monomial propagation.
@@
$f_{L}(\boldsymbol{\theta})=\langle H\rangle_{\boldsymbol{\theta}}={\rm Tr}\left[U_{L}(\boldsymbol{\theta})\varrho U_{L}(\boldsymbol{\theta})^{\dagger}H\right]$ (3)
given a unitary Fermionic circuit $U(\boldsymbol{\theta})$, an initial state $\varrho$, and an observable $H$. The initial state is typically a Fock basis state, $|\varrho=\outerproduct{n_{1}n_{2}\dots n_{N}}{n_{1}n_{2}\dots n_{N}}$. We further suppose the Fermionic circuit,
$U_{L}(\boldsymbol{\theta})=\prod_{j=1}^{L}e^{-\mathrm{i}\theta_{j}M_{{\boldsymbol{b}_{j}}}/2}\,$, (4)
@@
If the rotation angles for the circuit one wishes to simulate are known in advance then the sine and cosine terms take explicit numerical values and the $c_{\boldsymbol{b}}\left(\boldsymbol{\theta}\right)$ coefficients can be stored numerically. However, if the values of $\boldsymbol{\theta}$ are not known in advance, or one is interested in computing the expectation value in Eq. (3) for a wide range of $\boldsymbol{\theta}$ values, one can instead symbolically represent the coefficients $c_{\boldsymbol{b}}\left(\boldsymbol{\theta}\right)$. Such surrogate models have a substantial time and memory overhead, but after this pre-processing step, enable rapid re-evaluation of the expectation value [6, 8, 12].
@@
In its raw form, the number of monomials in the final expression for the expectation value, Eq. (6), will blow up exponentially with the number of gates in the circuit that induce a splitting. We therefore employ truncation schemes to drastically reduce the computational resources. There are two main families of truncations that can be employed to make the simulation more efficient, namely 1) ones that truncate small (or likely small) coefficients $c_{\boldsymbol{b}}$ and 2) ones that truncate any monomial $M_{\boldsymbol{b}}$ where Tr[$\Tr[\varrho M_{\boldsymbol{b}}]$ is small (or likely small). Inevitably these truncations will reduce the accuracy of the simulation but in many cases, the reduction in accuracy can be kept small.
@@
The second truncation strategy, whereby one aims to cut terms from the sum in Eq. (6) where $\text{Tr}[\varrho M_{\boldsymbol{b}}]$ is likely small, is more subtle. While potentially there are a number of different strategies one could take here, we will make use of the observation that in order for a term $\text{Tr}[\varrho M_{\boldsymbol{b}}]$ to be non-zero all the Majorana operators in the monomial $M_{\boldsymbol{b}}$ must be paired. That is, $\text{Tr}[\varrho M_{\boldsymbol{b}}]\neq 0$ if and only if, for every even Majorana operator $m_{2i}$ in the Majorana monomial, $m_{2i+1}$ is also in the monomial (this can be seen immediately from the definition of the Majorana operator and noting that $\langle\varphi|a_{i}|\varphi\rangle=\langle\varphi|a_{i}^{\dagger}|\varphi\rangle=0$ for any Fock basis state $|\varphi\rangle$). Next we make observations that for a randomly chosen Majorana monomial the higher its length $w$ the lower the probability that the monomial is paired (for $w<N$). This motivates a length truncation strategy. At each step, after each gate is applied, all terms corresponding to Majorana monomials with length above a certain threshold are to be discarded. This idea is depicted in Fig. 1.
While high length monomials are likely to not contribute to a Fock state this fact alone does not justify Majorana length truncation. Namely, high-length Majorana monomials could flow back into the low-length subspace, so eliminating them along the way could introduce significant errors. In broad strokes, the intuition behind why this does not cause issues is that 1) monomials tend to increase in length throughout the evolution, stabilizing at lengths $\approx N$ where their contribution to the estimation is irrelevant and 2) any monomials that do back flow tend to have small coefficients (due to the conservation of the 2-norm of the operator). We discuss these intuitions in more detail in the end matter and formalize them into error guarantees in the following section.
Error Guarantees.
Here we present guarantees for simulating typical Majorana circuits using low-length Majorana propagation. Concretely, we consider the Majorana circuit defined in Eq. (4) of depth $L$ and suppose that each Majorana generator $M_{{\boldsymbol{b}_{j}}}$ is a randomly chosen length $k=4$ monomial and each rotation angle is randomly chosen in the range $0,2\pi]$. This class of circuits can be regarded as a mathematically tractable, oversimplified model of certain relevant circuits, such as iteratively constructed variational circuits [53].
For simplicity in our analysis below we will suppose that the initial observable is a randomly chosen homogeneous monomial. By homogeneous we mean that the individual Majorana coefficients are unbiased, uncorrelated and with a variance that only depends on length - for a more formal definition see Def. 7 in the Appendices. We believe this is a reasonable toy model for a randomly chosen molecular Hamiltonian. The theorem can also be applied more widely to certain classes of biased distributions e.g. a random Majorana mode $O$ of length $o$.
We denote the approximation of $f_{L}(\boldsymbol{\theta})$ obtained by truncating all Majorana operators with length greater than $w_{0}$ by $f_{L}^{(w_{0})}(\boldsymbol{\theta})$. The average approximation error can quantified via the mean squared error (MSE),
$\Delta(w_{0},L)=\operatorname{\mathbb{E}}((f_{L}(\boldsymbol{\theta})-f_{L}^{(w_{0})}(\boldsymbol{\theta}))^{2})\,.$ (7)
This error can be bounded as follows.
Theorem 1 (Error analysis).
Consider a random Majorana circuit as described above, with a Fock basis initial state and a random homogeneous observable $O$. Assume that $w_{0}\geqslant 2$. Then for circuits with $L\leqslant 2N\ln(5ew_{0}4^{w_{0}})$:
$\frac{\Delta(w_{0},L)}{\operatorname{\mathbb{E}}\left(\|O\|_{2}^{2}\right)}\leqslant\frac{1}{2^{N-1}}+\left(\frac{ew_{0}}{N}\right)^{w_{0}/2}.$ (8)
Otherwise, for deeper circuits ($L\geqslant 2N\ln(5ew_{0}4^{w_{0}})$):
$\frac{\Delta(w_{0},L)}{\operatorname{\mathbb{E}}\left(\|O\|_{2}^{2}\right)}\leqslant\frac{1}{2^{N-1}}+5ew_{0}\,\left(\frac{16ew_{0}}{N}\right)^{w_{0}/2}e^{-L/2N}.$ (9)
[…]
Theorem 1 provides rigorous guarantees on the truncation error incurred when simulating Majorana circuits using a low-length approximation. As illustrated in Fig. 2, in all cases the error decreases exponentially with the truncation length $w_{0}$, reflecting that circuits can be faithfully captured using few-body Majorana operators. The comparison between the exact error (solid blue) and the theoretical bound (dashed green) shows that the empirical scaling closely follows the predicted behavior, validating the theoretical estimates.
Figure 2: Majorana Propagation error analysis. Comparison between the exact mean squared simulation error (computed via Theorem 5 in the end matter) and the theoretical error bound for a system of $N=100$ Majorana modes and initial observable Majorana length =4=4 for different length truncations $w_{0}$. The solid blue line shows the exact error as a function of circuit depth $L$, while the dashed green line represents the analytical upper bound on this error. The inset zooms in on low depths ($L=0$ to $L=100$). Here the observable $O$ is a random Majorana mode of length 4.
We note that for deeper circuits, in addition to the exponential suppression of the error with the cut off $w_{0}$, there is also an exponential suppression in circuit depth $L$. This is a reflection of the fact that the absolute value of expectation values of random circuits decay exponentially with circuit depth to a saturation value that is itself exponentially small in $N$ (a phenomenon analogous to barren plateaus from variational circuits with fixed ansatze [55, 56]). It is satisfying that our error bound, in contrast to prior work [10] in the context of Pauli propagation, carefully tracks this depth dependence.
We further bound the time complexity of our algorithm using the fact that the total number of Majorana monomials with length at most $w_{0}$ is $\mathcal{O}((2N)^{w_{0}})$. We can then combine Theorem 1 with Markov’s inequality to transform the average error statement into a probabilistic statement. On doing so, we obtain the following Theorem.
Theorem 2 (Time complexity).
Let $U$ be a randomly sampled circuit from an $L$-layered random Majorana rotation circuit on $2N$ modes, and let $O$ be a random homogeneous observable. Moreover, let $\epsilon,\delta>0$ such that $\epsilon^{-1}\delta^{-1}\leqslant 2^{o(N)}$. There exists a truncation length $w_{0}$ for which Majorana propagation runs in time $L\cdot N^{\mathcal{O}\left(\log(\epsilon^{-1}\delta^{-1})\right)},$ and outputs a value $f_{L}^{(w_{0})}(\vec{\theta})$, such that
[…]
for at least $1-\delta\mathbb{E}(\|O\|_{2}^{2})$ fraction of the circuits, if $N$ is large enough.
Theorem 2 thus establishes that for any constant error, only polynomial resources are required to classically simulate the circuit.
Implementation.
Figure 3: Majorana Propagation implementation for ground states. Top: The blue curve depicts the error in the estimated energy as a function of monomial length cut-off $w^{\star}$ for a 28-mode MP-simulated ADAPT-VQE circuit containing 22 Fermionic gates, computed by comparison against statevector simulation (SV). The error of the MP simulation decreases exponentially from below chemical precision (1.6 mHa) for $w^{\star}=4$ to around numerical precision for $w^{\star}=10$. The green curve shows the same quantity for 5 circuits obtained by uniformly randomizing angles in the range [0.1,0.1][-0.1,0.1]. The shaded area indicates the maximal and minimal errors attained. Bottom: The ADAPT-VQE inspired algorithm is run classically using MP to simulate the circuits for increasing problem sizes (solid lines). The $x$-axis reflects the number of Fermionic gates in the circuit. The $y$-axis value is the error in the energy, computed with MP with $w^{\star}=6$, divided by the spectral gap of the Hamiltonian, which bounds the lack of overlap with the ground state. For 28 modes we verify the accuracy of the MP simulation by computing the equivalent error via a SV simulation (yellow markers). SV simulation for the larger problems was not tractable. Only every 5th marker is included to improve readability.
Number Majorana Propagation Qiskit CUDA-Q CUDA-Q
of modes pre-pro. estim. SV SV MPS
28 0.05 s 0.0009 s 2.72 h 12.18 min 32.15 min
40 1.34 s 0.0395 s 10.10 h
52 13.85 s 0.1092 s >> 24 h
In Fig. 3 (top), we plot the error in the energy estimated with MP as a function of the length cut-off $w^{\star}$ for an ADAPT-VQE circuit comprising 22 Fermionic gates—equivalent to a 616-CNOT circuit on all-to-all connectivity (see Appendix E)—optimized for the 28-mode active space Hamiltonian. The error $\varepsilon$ is computed as $\varepsilon=|E_{\rm MP}(w^{\star})-E_{\rm sv}|$, where $E_{\rm sv}$ is the exact energy (up to numerical precision) obtained via statevector simulation. Remarkably, even for $w^{\star}=4$, MP produces an estimate with an error below chemical precision (1.6 mHa), with the error decreasing approximately exponentially as a function of the cut-off to nearly numerical precision for $w^{\star}=10$. A similar behaviour is observed when the angles are randomized within the parameter range in the optimized circuit, although the decay in $w^{\star}$ is slower.
In many applications, the relevant figure of merit quantifying the quality of a reference state $|\varphi\rangle$ is its overlap with the ground state $|\langle\varphi|\psi_{0}\rangle|$. Given the error in the energy, $|E-E_{0}|$, the overlap can be lower-bounded if the spectral gap of the Hamiltonian is known (see Appendix F). Thus, using the spectral gap computed with DMRG, we can upper-bound the state error, which we define as $1-|\langle\varphi|\psi_{0}\rangle|$. The bottom plot in Fig. 3 shows the upper bound to the state error as a function of the number of Fermionic gates in the circuit as the adaptive training progresses. All cases show a rapid convergence, reaching an overlap of at least 97% with the ground state with 22, 102, and 300 Fermionic gates for the three active space sizes (the simulations were manually interrupted upon reaching a absolute error of 10 mHa). When mapped to qubit space and transpiled for an all-to-all connectivity quantum computer, the equivalent circuits contain 616, 3470, and 13,676 CNOTs, respectively, which highlights the complexity of the obtained states.
Finally, Table 1 reports run-time benchmarks for energy estimation with MP and other methods: two statevector simulators (Qiskit and CUDA-Q), as well as CUDA-Q MPS simulator, a tensor network-based circuit simulator. The circuits used are the converged ADAPT-VQE circuits from Fig. 3 (bottom), transpiled to qubit space, resulting in the CNOT counts reported above. While statevector simulators are exact (up to numerical precision), they scale exponentially. Thus, only the 28-mode circuit could be simulated with these simulators with reasonable resources. MP and CUDA-Q MPS, on the other hand, produce approximate estimates but run in polynomial time. Details on the implementation of these methods can be found in Appendix E. Interestingly, in addition to simulating these circuits considerably faster than CUDA-Q MPS, the 28-mode result is also nearly two orders of magnitude more accurate, approximating the statevector energy within $\approx 1.39\cdot 10^{-2}$ mHa, as opposed to the $\approx 2.71$ mHa error in the MPS simulation.
Discussion.
Our analytical results portray MP as a very natural framework for the simulation of Fermionic circuits. MP leverages newly identified phenomena in the behaviour of basis monomials, namely monomial length, which is conceptually very different to Pauli weight. The underlying assumption in our analysis, specifically that circuits are unstructured, with gate generators randomly chosen from a set of bounded-length monomials, seems particularly appropriate to model iteratively constructed ground state approximating circuits. Our analysis also sheds light on the working principles behind approaches like Majorana-length-damping in tensor network methods for Fermionic systems [62].
Our numerical results substantiate that MP can be used to classically simulate ADAPT-VQE to produce circuits with significant overlap with the ground state in strongly correlated systems. In our simulations, MP was able to outperform state-of-the-art, tensor network circuit simulators. It would be interesting to compare its performance to more recently proposed approximate simulation methods for Fermionic circuits, like Refs. [63, 64, 65, 66]. Since MP does not rely on the sparsity of the wave-function in the Fock basis, and since single-excitation gates do not require additional truncation in MP, MP may produce more accurate estimations for circuits with many such excitations leading to dense states. Similarly, we stress that MP is strictly more powerful and efficient than ‘gsim’ approaches [67, 68, 69, 70, 71] as the basis of the algebra is effectively generated ‘on the fly’ during the simulation (rather than in advance) and so is not restricted to free Fermionic models with polynomially scaling Lie algebras [71, 72].
End Matter.